OpenStudy (anonymous):

consider a right triangle with legs of length a and b and hypotenuse of length c if k denotes the area of the triangle then value of(a+b) equals

OpenStudy (anonymous):

Area of a triangle is half the base times the height so: 0.5 x a x b = k Hypotenuse we find using pythagaras: a^2 + b^2 = c^2 c = sqrt(a^2+b^2) Not sure what you mean by the value of (a+b). In terms of what?

OpenStudy (triciaal):